# GeoGen: A Two-stage Coarse-to-Fine Framework for Fine-grained Synthetic Location-based Social Network Trajectory Generation

### v1

@@ Introduction
Fine-grained location-based social network (LBSN) check-in trajectory data are essential for many real-world applications such as next Point of Interest (POI) recommendation (Jiang et al. 2024), urban mobility understanding (Wang et al. 2019), business location selection (Xie et al. 2016), advertising (Jeon et al. 2021), and pandemic intervention (Hao et al. 2020). However, accessing large-scale LBSN data has become increasingly challenging due to high collection/purchasing costs and growing privacy concerns. Therefore, to support large-scale LBSN data access, it is crucial to develop methods that can generate synthetic LBSN check-in trajectories that maintain high fidelity and utility while safeguarding user privacy.
To date, some models have been proposed (Isaacman et al. 2012; Jiang et al. 2016; Yin et al. 2017; Ouyang et al. 2018; Zhu et al. 2023b) to generate LBSN check-in trajectory data. For instance, Yin et al. (Yin et al. 2017) and Jiang et al. (Jiang et al. 2016) employed physics-based models like hidden Markov models to synthesize activity trajectories. Generative adversarial networks (GANs) (Ouyang et al. 2018; Feng et al. 2020; Yu et al. 2017; Yuan et al. 2022) have demonstrated substantial advances in generating LBSN check-in trajectories. However, these approaches inevitably suffer from an unstable and extremely long training procedure. The impressive achievements of diffusion models in various tasks, such as image (Dhariwal and Nichol 2021; Kulikov et al. 2023) and audio (Huang et al. 2022) generation, have sparked interest in their applications to trajectory generation. Several diffusion-based models (Zhu et al. 2023b, 2024; Liu et al. 2023) have been proposed to generate spatially continuous, temporally regular GPS trajectories with fixed intervals between consecutive data points and fixed lengths (e.g., 100 data points per trajectory). These models have demonstrated strong performance, highlighting the great potential of diffusion models for generating LBSN check-in trajectories.
@@
However, it is nontrivial to adapt diffusion models to LBSN check-in trajectory data generation due to its two unique natures: (i) Temporally irregular: As shown in Figure 1, in contrast to regularly sampled GPS trajectories, LBSN check-in trajectories exhibit complex spatio-temporal patterns due to uncertain human mobility and check-in behavior, leading to variable time intervals between points and unfixed trajectory lengths. (ii) Spatially discrete: Diffusion models are originally designed to generate fixed-length data in the continuous domain, while POIs are from a discrete spatial space, so it is challenging to adapt existing diffusion models designed for regularly-sampled trajectory generation (Zhu et al. 2023b, 2024; Liu et al. 2023) to LBSN check-in trajectories. In addition, a large time interval setting (e.g., 60 minutes) may decrease the accuracy for LBSN check-in trajectory generation, while a small interval (e.g., 1 minute) may significantly decrease the generation efficiency.
@@
To address these challenges, we propose GeoGen, a two-stage, coarse-to-fine framework for synthetic LBSN trajectory data generation. In the first stage, to adapt diffusion models to irregularly-sampled LBSN trajectory data, we reconstruct spatially continuous and temporally regular latent movement sequences from LBSN check-in trajectories through interpolation with a pre-determined coarse-grained interval. We then design a Sparsity-aware Spatio-temporal Diffusion model (S2TDiff), which includes a novel and efficient Spatially-Aware Sparsely-Gated U-Net (SASG-UNet) as its denoising network. SASG-UNet replaces standard convolutions with hierarchical 1D blocks to capture multi-scale patterns and integrates a specialized S2G Attention module to address the unique characteristics of the latent movement sequences. In the second stage, we propose a Transformer-based Seq2Seq architecture Coarse2FineNet that auto-regressively generates fine-grained check-in points based on the coarse-grained latent movement sequences. Coarse2FineNet incorporates a POI context-aware encoder that captures rich contextual representations of the latent movement sequence, and a Multi-task Hybrid-Head Decoder that generates the next POI and its fine-grained timestamp. In the encoder, we introduce a dynamic context fusion mechanism that aligns the input latent movement sequence with POIs exhibiting diverse spatio-temporal characteristics, integrating them into unified contextual representations. In the decoder, instead of directly predicting the next check-in point, we leverage a temporal point process to model the user behavior uncertainty underlying LBSN check-in trajectories. This two-stage approach effectively resolves a key trade-off between computational efficiency and data quality, enabling the generation of large-scale synthetic LBSN check-in trajectories.
@@
  • Technically, we propose a two-stage coarse-to-fine framework called GeoGen for synthetic LBSN trajectory generation. In the first stage, we design a novel S2TDiff model with an efficient SASG-UNet as a denoising network to generate coarse-grained latent movement sequences. In the second stage, we design a Transformer-based Seq2Seq architecture, Coarse2FineNet, which includes a POI context-aware encoder and a multi-task hybrid-head decoder to generate fine-grained POI visits with precise, non-uniform timestamps.
@@
Definition 1. (LBSN Check-in Trajectories). An LBSN check-in trajectory is described as a POI visit sequence $s=[x_{1},x_{2},\ldots]$ of a user. The $i$th element $x_{i}$ is a check-in point defined as a tuple $(p_{i},t_{i})$, where $p_{i}$ denotes a unique POI location ID corresponding to a geographic coordinate $g_{p_{i}}$, and $t_{i}$ is the check-in timestamp.
@@
Definition 3. (LBSN Check-in Trajectory Generation). Given a real-world dataset containing $m$ fine-grained LBSN check-in trajectories, denoted as $S=[s_{1},s_{2},\ldots]$, the LBSN check-in trajectory generation task aims to generate a synthetic trajectory dataset $\hat{S}=[\hat{s}_{1},\hat{s}_{2},\ldots]$ that maintains the spatio-temporal characteristics and distributions of $S$, as well as high data utility for downstream applications.
Figure 2: The overall architecture of our proposed GeoGen framework consists of two main stages. In the first stage, within S2TDiff, each reconstructed latent movement sequence $s_{\text{lat}}$ is first transformed into pure noise $s_{\text{lat}}^{N}$. During inference, SASG-UNet progressively denoises a sampled noise sequence to generate a synthetic latent trajectory. In the second stage, the context-aware encoder of Coarse2FineNet extracts contextualized representations from the generated latent movement sequence. The decoder then auto-regressively generates each POI and fine-grained timestamp pair, conditioned on both the contextual representations and the partially generated LBSN check-in trajectory. FFN and TPP refer to the feed-forward network and temporal point process, respectively.
@@
Denoising Diffusion Probabilistic Model (Ho, Jain, and Abbeel 2020) encompass two processes: (i) a forward diffusion process that gradually introduces noise into the data, and (ii) a backward diffusion process that reconstructs the original data from its noisy state. The forward diffusion process is a non-trainable Markov chain that transforms real sample $\mathbf{x}^{0}$ into latent variables $\mathbf{x}^{1},\ldots,\mathbf{x}^{N}$, which can be represented as:
@@
where $\beta_{1},\ldots,\beta_{N}\in(0,1)$ is a pre-determined variance schedule. $\mathbf{x}^{n}$ can be directly sampled from $\mathbf{x}^{0}$ with $q(\mathbf{x}^{n}|\mathbf{x}^{0})=\mathcal{N}(\mathbf{x}^{n};\sqrt{\overline{\alpha}_{n}}\mathbf{x}^{0},(1-\overline{\alpha}_{n})\mathbf{I})$, where $\alpha_{n}=1-\beta_{n}$ and $\overline{\alpha}_{n}=\prod_{n=1}^{N}\alpha_{n}$. With the application of the reparameterization, $\mathbf{x}^{n}$ can be represented as $\mathbf{x}^{n}=\sqrt{\overline{\alpha}_{n}}\mathbf{x}^{0}+\sqrt{1-\overline{\alpha}_{n}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. The backward diffusion process is a trainable Markov chain that aims to recover $\mathbf{x}^{0}$ from $\mathbf{x}^{N}$, which can be formulated as:
$p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})=\mathcal{N}(\mathbf{x}^{n-1};\mu_{\theta}(\mathbf{x}^{n},n),\sigma_{\theta}(\mathbf{x}^{n},n)\mathbf{I}),$ (2)
@@
$\mathbb{E}_{\mathbf{x}^{0},\boldsymbol{\epsilon}}\left[\frac{\beta_{n}^{2}}{2\Sigma_{\theta}\alpha_{n}(1-\overline{\alpha}_{n})}\left\|\boldsymbol{\epsilon}-\epsilon_{\theta}\left(\sqrt{\overline{\alpha}_{n}}\mathbf{x}^{0}+\sqrt{1-\overline{\alpha}_{n}}\boldsymbol{\epsilon},n\right)\right\|^{2}\right],$ (3)
where $\epsilon_{\theta}$ is a neural network for predicting sampled $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. After trained, trajectory generation is conducted by progressively sampling $\mathbf{x}^{n-1}$ from distribution $p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})$ until reach $\mathbf{x}^{0}$ by computing:
$\mathbf{x}^{n-1}=\frac{1}{\sqrt{\alpha_{n}}}\left(\mathbf{x}^{n}-\frac{\beta_{n}}{\sqrt{1-\alpha_{n}}}\epsilon_{\theta}(\mathbf{x}^{n},n)\right)+\sqrt{\Sigma_{\theta}}\mathbf{z},$ (4)
@@
To adapt diffusion models to irregularly-sampled LBSN trajectory data, we first convert original fine-grained LBSN trajectories into coarse-grained spatially continuous and temporally regular latent movement sequences. We then design a Sparsity-aware Spatio-temporal Diffusion model (S2TDiff) to generate coarse-grained latent movement sequences by learning underlying spatio-temporal distributions and physical constraints.
@@
We first reconstruct each real LBSN check-in trajectory $s$ into a latent movement sequence, $s_{\mathit{lat}}=[(l_{1},c_{1}),\ldots,(l_{L},c_{L})]$. Given a time interval $\mathit{I}$, we divide the total duration into $L$ slots, where $L=[D/I]$, $D$ is the duration. For each slot $i$, the latent coordinate $l_{i}$ is the geographic mean of all check-ins $P_{i}$ falling within that time window ($l_{i}=\frac{1}{|P_{i}|}\sum_{(p_{j},t_{j})\in P_{i}}g_{p_{j}}$), and the intensity $c_{i}$ is their count ($c_{i}=|P_{i}|$). Next, we adopt linear interpolation between the nearest preceding ($i_{p}$) and subsequent ($i_{n}$) time slots with no check-ins ($c_{i}=0$), where the missing location is calculated as $l_{i}=l_{i_{p}}+((i-i_{p})/(i_{n}-i_{p}))\cdot(l_{i_{n}}-l_{i_{p}})$. For empty slots at the sequence boundaries, we apply circular interpolation, treating the trajectory as a loop to infer locations between the first and last observed points. This reformulation transforms sparse event data into a spatially continuous and temporally regular format, enabling compatibility with diffusion models while reducing memory and computational costs via temporal compression.
@@
To capture the underlying spatio-temporal patterns, we propose a Sparsity-aware Spatio-temporal Diffusion model (S2TDiff), trained using latent movement sequences $S_{\text{lat}}=[s_{\text{lat},1},s_{\text{lat},2},\dots]$ as input. While interpolation yields spatially continuous sequences, it poses challenges of (1) diverse movement rates and (2) sparse intensity.
To address these challenges, we design a novel and efficient Spatially-Aware Sparsely-Gated U-Net (SASG-UNet) as the denoising backbone in S2TDiff. S2TDiff employs a U-Net backbone with hierarchical 1D convolution blocks that replace standard 2D convolutions. These blocks efficiently capture multi-resolution features using parallel branches of multi-scale average pooling and fixed-size 1D convolutions. At training/inference step $n$, the denoising network SASG-UNet $\epsilon_{\theta}$ takes a noised latent movement sequence as input and outputs the predicted noise: $\epsilon_{\theta}(s_{\text{lat}}^{n},n)$. To further tackle sparse intensity and diverse movement rates, we introduce a novel S2G Attention module within SASG-UNet’s skip connections, consisting of two complementary pathways: the Intensity-Gated Pathway and the Spatially-Aware Attention Pathway. Firstly, the Intensity-Gated Pathway handles sparse signals by with a lightweight convolutional sub-network to generate a dynamic gating signal $g=\sigma(\text{Conv1D}(\text{ReLU}(\text{Conv1D}(h_{\text{enc}}))))$, where $h_{\text{enc}}$ denotes encoder features and $\sigma$ is a sigmoid function. This gate adaptively modulates the features, allowing the network to prioritize learning from informative, sparse, non-zero intensity values. Secondly, the Spatially-Aware Attention Pathway accounts for diverse movement rates via a Local Spatial Bias $b$, where at index $i$, $b_{i}=\mathcal{F}_{\text{emb}}(d_{\text{H}}((s^{n}_{\text{lat},i},s^{n}_{\text{lat},i-1}))$, and $d_{\text{H}}$ is the Haversine distance (Maria et al. 2020) between consecutive points $s^{n}_{\text{lat},i}$, $s^{n}_{\text{lat},i-1}$ in the noisy input and $\mathcal{F}_{\text{emb}}$ is a learnable feed-forward network. Next, we add an attention mechanism, yielding the final refined features: $h_{\text{att}}=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$, where $Q$ and $K$ are the self-attention output features of the Intensity-Gated Pathway and $V$ is the bias $b$. This mechanism provides the model with a direct and efficient understanding of real-world movement dynamics, enabling it to handle sparse, irregular mobility patterns effectively.
@@
To reduce complexity in the diffusion process, we compress the latent sequence $s_{\text{latent}}=\{(l_{i},c_{i})\}$ into $s_{\text{latent}}^{\prime}=\{(l^{\prime}_{i},t^{\prime}_{i})\mid c_{i}\geq\gamma\}$ by filtering points, where $l^{\prime}_{j}=l_{i_{j}}$ and $t^{\prime}_{j}=t_{i_{j}}=(i_{j}-0.5)\cdot\mathit{Int}$, $\gamma$ denotes intensity threshold, and $t_{i}$ denotes the midpoint time of slot $i$. We first obtain spatial embeddings $l_{vec}=\{Linear(l^{\prime}_{i})\}$ and temporal embedding $t_{vec}=\{T2V(t^{\prime}_{i})\}$, where $T2V$ denotes the Time2Vec method  (Kazemi et al. 2019) that captures both linear and cyclical temporal patterns.
To establish local POI context awareness, we design a Dynamic Context Fusion (DCF) module within the encoder. In DCF, a shared POI embedding codebook is built to encode multiple attributes, including geographic coordinates $p_{j}$, visit frequency vectors $f_{j}$, and category one-hot vector $c_{j}$ for POI $j$: $e_{\text{POI}}^{(j)}=W_{\text{latlon}}p_{j}+W_{\text{freq}}f_{j}+W_{\text{cat}}c_{j}$, where $W_{\text{latlon}},W_{\text{freq}},W_{\text{cat}}$ are learnable parameters. Then, we employ a dual attention mechanism to model spatial and temporal relationships separately. The spatial attention maps each coordinate $l^{\prime}_{i}$ from the filtered sequence to a distribution over POIs based on proximity. The attention weight $\alpha_{j}$ for each POI $j$ and coordinate $l_{i}$ is calculated as:
@@
where $\tau$ is a temperature parameter that controls the distribution’s sharpness and $g_{p_{j}}$ denotes the geographic coordinate of POI $p_{j}$. Concurrently, the temporal attention captures time-specific visit patterns. The temporal embedding $t^{\prime}_{vec,i}$ is projected into a query vector $w_{i}=W_{t}t^{\prime}_{vec,i}$, and the attention weight is computed via dot-product similarity:
$\alpha_{i,j}^{(t)}=\frac{\exp(f_{j}\cdot w_{i})}{\sum_{k=1}^{|\mathcal{P}|}\exp(f_{k}\cdot w_{i})}$ (6)
Finally, we adaptively balance the spatial and temporal signals by computing fusion weights using a softmax over their average representations $(\beta_{s},\beta_{t})=\text{Softmax}\left(\mathbf{W}_{f}\left[\frac{1}{L}\sum_{i=1}^{L}\mathbf{l}_{i}\,\|\,\frac{1}{L}\sum_{i=1}^{L}\mathbf{t}_{i}\right]\right)$ , where $\mathbf{l}_{i}$ and $\mathbf{t}_{i}$ denote the $i$-th spatial and temporal feature vectors, respectively; $\mathbf{W}_{f}$ is a learnable linear projection matrix that is used to create a unified attention distribution $\boldsymbol{\alpha}_{i,j}=\beta_{s}\boldsymbol{\alpha}^{(s)}_{i,j}+\beta_{t}\boldsymbol{\alpha}^{(t)}_{i,j}$. This fused attention is applied to the shared POI embeddings to create a context-aware representation $h_{\text{poi},i}=\sum_{j=1}^{|\mathcal{P}|}\alpha_{i,j}e_{\text{POI}}^{(j)}$. This POI-enhanced vector $h_{poi}$ is integrated with the initial features $h_{vec}=[l_{vec},t_{vec}]$ through a residual connection, and is processed by a Transformer Encoder to produce the contextualized representations $H$.
Table 1: Fidelity Evaluation. The best results on each dataset are in bold, and the second-best results are underlined.
Method Metric Distance Radius Interval Length Distance Radius Interval Length

### v2

@@ Introduction
Fine-grained location-based social network (LBSN) check-in trajectories are essential for many real-world applications such as next Point of Interest (POI) recommendation (Jiang et al. 2024), urban mobility understanding (Wang et al. 2019), business location selection (Xie et al. 2016), advertising (Jeon et al. 2021), and pandemic intervention (Hao et al. 2020). However, accessing large-scale LBSN data has become increasingly challenging due to high collection/purchasing costs and growing privacy concerns. Therefore, to support large-scale LBSN data access, it is crucial to develop methods that can generate synthetic LBSN check-in trajectories that maintain high fidelity and utility while safeguarding user privacy.
@@
To date, some methods have been proposed (Isaacman et al. 2012; Jiang et al. 2016; Yin et al. 2017; Ouyang et al. 2018; Zhu et al. 2023b) to generate LBSN check-in trajectory data. For instance, Generative adversarial networks (GANs) (Ouyang et al. 2018; Feng et al. 2020; Yu et al. 2017; Yuan et al. 2022) have demonstrated substantial advances in generating LBSN check-in trajectories. However, these approaches inevitably suffer from an unstable and extremely long training procedure. Recently, the successes of diffusion models across modalities, including image (Dhariwal and Nichol 2021; Kulikov et al. 2023) and audio (Huang et al. 2022) generation, have spurred investigations into trajectory generation. Several diffusion-based models (Zhu et al. 2023b, 2024; Liu et al. 2023) have been proposed to generate spatially continuous, temporally regular GPS trajectories with fixed intervals between consecutive data points and fixed lengths (e.g., 100 data points per trajectory). These models have demonstrated strong performance, highlighting the great potential of diffusion models for generating LBSN check-in trajectories.
@@
However, it is nontrivial to adapt diffusion models to LBSN check-in trajectory data generation due to its two unique natures: (i) Temporally irregular: As shown in Figure 1, in contrast to regularly sampled GPS trajectories, LBSN check-in trajectories exhibit complex spatio-temporal patterns due to uncertain human mobility and check-in behavior, leading to variable time intervals between points and unfixed trajectory lengths. (ii) Spatially discrete: Diffusion models are well suited to fixed-length data in the continuous domain, whereas POIs lie in a discrete spatial space. Consequently, adapting models designed for regularly sampled trajectory generation (Zhu et al. 2023b, 2024; Liu et al. 2023) to LBSN check-in trajectories is challenging: a large sampling interval (e.g., 60 minutes) degrades fidelity, while a small interval (e.g., 1 minute) severely reduces efficiency.
@@
To address these challenges, we propose GeoGen, a two-stage, coarse-to-fine framework for synthetic LBSN trajectory data generation. In the first stage, to adapt diffusion models to irregularly-sampled LBSN trajectory data, we reconstruct spatially continuous and temporally regular latent movement sequences from LBSN check-in trajectories through interpolation with a pre-determined coarse-grained interval. We then design a Sparsity-aware Spatio-temporal Diffusion model (S2TDiff), which includes a novel and efficient Spatially-Aware Sparsely-Gated U-Net (SASG-UNet) as its denoising network. SASG-UNet replaces standard convolutions with hierarchical 1D blocks to capture multi-scale patterns and integrates a specialized S2G Attention module to fit the unique characteristics of the latent movement sequences. In the second stage, we propose a Transformer-based Seq2Seq architecture Coarse2FineNet that auto-regressively generates fine-grained check-in points based on the coarse-grained latent movement sequences. Coarse2FineNet incorporates a POI Context-aware Encoder that captures rich contextual representations of the latent movement sequence, and a Multi-task Hybrid-Head Decoder that generates the next POI and its fine-grained timestamp. In the encoder, we introduce a dynamic context fusion mechanism that aligns the input latent movement sequence with POIs exhibiting diverse spatio-temporal characteristics, integrating them into unified contextual representations. In the decoder, instead of directly predicting the next check-in point, we leverage a temporal point process to model the user behavior uncertainty underlying LBSN check-in trajectories. This two-stage approach effectively resolves a key trade-off between computational efficiency and data quality, enabling the generation of large-scale synthetic LBSN check-in trajectories.
@@
  • Technically, we propose a two-stage coarse-to-fine framework called GeoGen for synthetic LBSN trajectory generation. In the first stage, we design a novel S2TDiff model with an efficient SASG-UNet as a denoising network to generate coarse-grained latent movement sequences. In the second stage, we design a Transformer-based Seq2Seq architecture, Coarse2FineNet, which includes a POI Context-aware Encoder and a Multi-task Hybrid-head Decoder to generate fine-grained check-ins at precise, non-uniform timestamps.
@@
Definition 1. (LBSN Check-in Trajectories). An LBSN check-in trajectory is a sequence of POI visits from an individual, denoted $s=[x_{1},x_{2},\ldots]$. The $i$th element $x_{i}$ is a check-in represented as the tuple $x_{i}=(p_{i},t_{i})$, where $p_{i}\in P$ is a POI ID from a local finite set $P$ with associated geographic coordinate $g_{p_{i}}$, and $t_{i}$ is the check-in timestamp.
@@
Definition 3. (LBSN Check-in Trajectory Generation). Given a real-world dataset containing fine-grained LBSN check-in trajectories, denoted as $S=[s_{1},s_{2},\ldots]$, the LBSN check-in trajectory generation task aims to generate a synthetic trajectory dataset $\hat{S}=[\hat{s}_{1},\hat{s}_{2},\ldots]$ that maintains the spatio-temporal characteristics and distributions of $S$, as well as high data utility for downstream applications.
Figure 2: The overall framework of our proposed GeoGen. In the first stage, within S2TDiff, each reconstructed latent movement sequence $s_{\text{lat}}$ is first transformed into pure noise $s_{\text{lat}}^{N}$. During inference, SASG-UNet progressively denoises a sampled noise sequence to generate a synthetic latent trajectory. In the second stage, the context-aware encoder of Coarse2FineNet extracts contextualized representations from the generated latent movement sequence. The decoder then auto-regressively generates each POI and fine-grained timestamp pair, conditioned on both the contextual representations and the partially generated LBSN check-in trajectory. FFN and TPP refer to the feed-forward network and temporal point process, respectively.
@@
Denoising Diffusion Probabilistic Models (Ho, Jain, and Abbeel 2020) encompass two processes: (i) a forward diffusion process that gradually introduces noise into the data, and (ii) a backward diffusion process that reconstructs the original data from its noisy state. The forward diffusion process is a non-trainable Markov chain that transforms real sample $\mathbf{x}^{0}$ into latent variables $\mathbf{x}^{1},\ldots,\mathbf{x}^{N}$, which can be represented as:
@@
where $\beta_{1},\ldots,\beta_{N}\in(0,1)$ are pre-determined variance schedules. $\mathbf{x}^{n}$ can be directly sampled from $\mathbf{x}^{0}$ with $q(\mathbf{x}^{n}|\mathbf{x}^{0})=\mathcal{N}(\mathbf{x}^{n};\sqrt{\overline{\alpha}_{n}}\mathbf{x}^{0},(1-\overline{\alpha}_{n})\mathbf{I})$, where $\alpha_{n}=1-\beta_{n}$ and $\overline{\alpha}_{n}=\prod_{i=1}^{N}\alpha_{i}$. With the application of the reparameterization, $\mathbf{x}^{n}$ can be represented as $\mathbf{x}^{n}=\sqrt{\overline{\alpha}_{n}}\mathbf{x}^{0}+\sqrt{1-\overline{\alpha}_{n}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. The backward diffusion process is a trainable Markov chain that aims to recover $\mathbf{x}^{0}$ from $\mathbf{x}^{N}$, which can be formulated as:
$p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})=\mathcal{N}(\mathbf{x}^{n-1};\mu_{\theta}(\mathbf{x}^{n},n),\sigma_{\theta}^{2}(\mathbf{x}^{n},n)\mathbf{I}),$ (2)
@@
$\mathbb{E}_{\mathbf{x}^{0},\boldsymbol{\epsilon},\,n}\!\left[\left\|\boldsymbol{\epsilon}-\epsilon_{\theta}\!\left(\sqrt{\overline{\alpha}_{n}}\mathbf{x}^{0}+\sqrt{1-\overline{\alpha}_{n}}\,\boldsymbol{\epsilon},\,n\right)\right\|^{2}\right].$ (3)
where $\epsilon_{\theta}$ is a neural network for predicting sampled $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. After training, trajectory generation is conducted by progressively sampling $\mathbf{x}^{n-1}$ from distribution $p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})$ until reach $\mathbf{x}^{0}$ by computing:
$\mathbf{x}^{n-1}=\frac{1}{\sqrt{\alpha_{n}}}\!\left(\mathbf{x}^{n}-\frac{\beta_{n}}{\sqrt{1-\overline{\alpha}_{n}}}\,\epsilon_{\theta}(\mathbf{x}^{n},n)\right)+\sigma_{\theta}(\mathbf{x}^{n},n)\,\mathbf{z},$ (4)
@@
To adapt diffusion models to irregularly-sampled LBSN trajectory data, we first convert original fine-grained LBSN trajectories into coarse-grained spatially continuous and temporally regular latent movement sequences. We then design a Sparsity-aware Spatio-temporal Diffusion model (S2TDiff) to generate synthetic coarse-grained latent movement sequences by learning underlying spatio-temporal distributions and physical constraints.
@@
Given a time interval $\mathit{I}$, we divide the total duration into $L$ slots, where $L=[D/I]$, $D$ is the duration. We reconstruct each real LBSN check-in trajectory $s$ into a latent movement sequence, $s_{\mathit{lat}}=[(l_{1},c_{1}),\ldots,(l_{L},c_{L})]$. For each slot $i$, the latent coordinate $l_{i}$ is the geographic mean of all check-ins $P_{i}$ falling within that time window ($l_{i}=\frac{1}{|P_{i}|}\sum_{(p_{j},t_{j})\in P_{i}}g_{p_{j}}$), and the intensity $c_{i}$ is their count ($c_{i}=|P_{i}|$). Next, we adopt linear interpolation between the nearest preceding ($i_{p}$) and subsequent ($i_{n}$) time slots with no check-ins ($c_{i}=0$), where the missing location is calculated as $l_{i}=l_{i_{p}}+((i-i_{p})/(i_{n}-i_{p}))\cdot(l_{i_{n}}-l_{i_{p}})$. For empty slots at the sequence boundaries, we apply circular interpolation, treating the trajectory as a loop to infer locations between the first and last observed points. This reformulation transforms sparse event data into a spatially continuous and temporally regular format, enabling compatibility with diffusion models while reducing memory and computational costs via temporal compression.
@@
To capture the underlying spatio-temporal patterns, we propose a Sparsity-aware Spatio-temporal Diffusion model (S2TDiff), trained with latent movement sequences $S_{\text{lat}}=[s_{\text{lat},1},s_{\text{lat},2},\dots]$ as input. Although linear interpolation yields spatially continuous sequences, it poses challenges of (1) diverse movement rates and (2) sparse intensity.
To address these challenges, we design a novel and efficient Spatially-Aware Sparsely-Gated U-Net (SASG-UNet) as the denoising network in S2TDiff. S2TDiff employs a U-Net backbone with hierarchical 1D convolution blocks that replace standard 2D convolutions. These blocks efficiently capture multi-resolution features using parallel branches of multi-scale average pooling and fixed-size 1D convolutions. At training/inference step $n$, the denoising network SASG-UNet $\epsilon_{\theta}$ takes a noisy latent movement sequence $s_{\text{lat}}^{n}$ as input and outputs the predicted noise: $\epsilon_{\theta}(s_{\text{lat}}^{n},n)$. To further tackle sparse intensity and diverse movement rates, we introduce a novel S2G Attention module within SASG-UNet’s skip connections, consisting of two complementary pathways: the Intensity-Gated Pathway and the Spatially-Aware Attention Pathway. Firstly, the Intensity-Gated Pathway handles sparse signals by with a lightweight convolutional sub-network to generate a dynamic gating signal $g=\sigma(\text{Conv1D}(\text{ReLU}(\text{Conv1D}(h_{\text{enc}}))))$, where $h_{\text{enc}}$ denotes encoder features and $\sigma$ is a sigmoid function. This gate adaptively modulates the features, allowing the network to prioritize learning from informative, sparse, non-zero intensity values. Secondly, the Spatially-Aware Attention Pathway accounts for diverse movement rates via a Local Spatial Bias $b$, where at index $i$, $b_{i}=\mathcal{F}_{\text{emb}}(d_{\text{H}}((s^{n}_{\text{lat},i},s^{n}_{\text{lat},i-1}))$, and $d_{\text{H}}$ is the Haversine distance (Maria et al. 2020) between consecutive points $s^{n}_{\text{lat},i}$, $s^{n}_{\text{lat},i-1}$ in the noisy input and $\mathcal{F}_{\text{emb}}$ is a learnable feed-forward network. Next, we add an attention mechanism, yielding the final refined features: $h_{\text{att}}=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$, where $Q$ and $K$ are the self-attention output features of the Intensity-Gated Pathway and $V$ is the bias $b$. This mechanism provides the model with a direct and efficient understanding of real-world movement dynamics, enabling it to handle sparse, irregular mobility patterns effectively.
@@
To reduce complexity in the diffusion process, we compress a latent sequence $s_{\text{latent}}=\{(l_{i},c_{i})\}$ into $s_{\text{latent}}^{\prime}=\{(l^{\prime}_{i},t^{\prime}_{i})\mid c_{i}\geq\gamma\}$ by filtering points, where $l^{\prime}_{j}=l_{i_{j}}$ and $t^{\prime}_{j}=t_{i_{j}}=(i_{j}+0.5)\cdot\mathit{I}$ , $\gamma$ denotes intensity threshold, and $t_{i}$ denotes the midpoint time of slot $i$. We first obtain spatial embeddings $l_{vec}=Linear(l^{\prime}_{i})$ and temporal embedding $t_{vec}=T2V(t^{\prime}_{i})$, where $T2V$ denotes the Time2Vec method  (Kazemi et al. 2019) that captures both linear and cyclical temporal patterns.
To establish local POI context awareness, we design a Dynamic Context Fusion (DCF) module within the encoder. In DCF, a shared POI embedding codebook is built to encode multiple attributes, including geographic coordinates $g_{p_{j}}$, temporal visit frequency vectors $f_{p_{j}}$, and category one-hot vector $c_{p_{j}}$ for POI $p_{j}$}: $e_{\text{POI}}^{(j)}=W_{\text{latlon}}g_{p_{j}}+W_{\text{freq}}f_{p_{j}}+W_{\text{cat}}c_{p_{j}}$, where $W_{\text{latlon}},W_{\text{freq}},W_{\text{cat}}$ are learnable weights. Then, we employ a dual attention mechanism to model spatial and temporal relationships separately. The spatial attention maps each coordinate $l^{\prime}_{i}$ from the filtered sequence to a distribution over POIs based on proximity. The attention weight $\alpha_{j}$ for each POI $j$ and coordinate $l_{i}$ is calculated as:
@@
where $\tau_{s}$ is a temperature parameter that controls the distribution’s sharpness. Concurrently, the temporal attention captures time-specific visit patterns. The temporal embedding $t^{\prime}_{vec,i}$ is projected into a query vector $w_{i}=W_{t}t^{\prime}_{vec,i}$, and the attention weight is computed via dot-product similarity:
$\alpha_{i,j}^{(t)}=\frac{\exp(f_{p_{j}}\cdot w_{i})}{\sum_{k=1}^{|\mathcal{P}|}\exp(f_{p_{k}}\cdot w_{i})}$ (6)
Finally, we adaptively balance the spatial and temporal signals by computing fusion weights using a softmax over their average representations $(\beta_{s},\beta_{t})=\text{Softmax}\left(\mathbf{W}_{\alpha}\left[\frac{1}{L}\sum_{i=1}^{L}\mathbf{l}_{i}\,\|\,\frac{1}{L}\sum_{i=1}^{L}\mathbf{t}_{i}\right]\right)$ , where $\mathbf{l}_{i}$ and $\mathbf{t}_{i}$ denote the $i$-th spatial and temporal feature vectors, respectively; $\mathbf{W}_{\alpha}$ is a learnable linear projection matrix that is used to create a unified attention distribution $\boldsymbol{\alpha}_{i,j}=\beta_{s}\boldsymbol{\alpha}^{(s)}_{i,j}+\beta_{t}\boldsymbol{\alpha}^{(t)}_{i,j}$. This fused attention is applied to the shared POI embeddings to create a context-aware representation $h_{\text{poi},i}=\sum_{j=1}^{|\mathcal{P}|}\alpha_{i,j}e_{\text{POI}}^{(j)}$. This POI-enhanced vector $h_{poi}$ is integrated with the initial features $h_{vec}=[l_{vec},t_{vec}]$ through a residual connection, and is processed by a Transformer Encoder to produce the contextualized representations $H$.
Method Metric Distance Radius Interval Length Average Distance Radius Interval Length Average